OpenStudy (anonymous):

How do I find the square root for the following equation? dr=-4sin(t)*dt+3*dt+4cos(t)*dt

OpenStudy (baldymcgee6):

what do you mean?

OpenStudy (anonymous):

How do I solve for dt

OpenStudy (anonymous):

\[\int\limits dr =\int\limits -4\sin(t)*dt+\int\limits 3*dt+\int\limits 4\cos(t)*dt\]

OpenStudy (anonymous):

4cos(t)+3x+4Sin(t)

OpenStudy (anonymous):

r = 4cos(t)+3t+4sin(t) + c

OpenStudy (anonymous):

why is my text getting sqrt13 as the answer?

OpenStudy (anonymous):

go you have value for t?

OpenStudy (anonymous):

0<=t<=3

OpenStudy (anonymous):

Anyone?